OpenStudy (anonymous):

It costs $900 to start up a business of selling hot dogs. Each hot dog costs $0.75 to produce. Let y be the cost in dollars of producing x hot dogs. Write the cost equation.

OpenStudy (anonymous):

1) y= 0.75x +900 Because every hot dog costs $0.75 you multiply 0.75 by 'x' (the number of hot dogs). Then you add 900 because that is how much you have to pay even if you don't make any hot dogs. 2) Slope intercept form is y=mx+b where y and x are the values of a given point (2, -7) , m is the slope, and b is the y- intercept. To find the slope you have to rearrange the equation -5x +3y = 11 into slope intercept form so that y is on one side. So you get y= (5/3)x+(11/3). The slope of this line is 5/3. You want the slope that is perpendicular to this line so you take the negative reciprocal. Which basically means flip the fraction and make it negative. So your new slope is (-3/5). Now you have y (-7), x (2), and, m(-3/5). Plug them into the equation y=mx+b to solve for b. y=mx+b -7=(-3/5)(2) +b b= -29/5 So your final answer is y= (-3/5)x -(29/5)